OpenStudy (anonymous):

Solution to 3x/4 + 1/4?

OpenStudy (anonymous):

\[\frac{ 3x }{ 4 }+\frac{ 1 }{ 4}=0\]

OpenStudy (anonymous):

=5, sorry

OpenStudy (anonymous):

okay what do you notice is common between the two fractions?

OpenStudy (anonymous):

the denominator :)

OpenStudy (anonymous):

so multiply both sides by 4 (the denominator)

OpenStudy (anonymous):

the expressions are equal to five, did you see that, just making sure

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@scorpions786 I need you :c

OpenStudy (anonymous):

OKay still multiply by 4.

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

I dont understand :c why multiply by four if they are already the same?

OpenStudy (anonymous):

You are trying to solve for x correct?

OpenStudy (anonymous):

yes :) is that why?

OpenStudy (anonymous):

let me show you what I mean...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what you have is this...\[\frac{ 3x+1 }{ 4}=5\]

OpenStudy (anonymous):

now you want to isolate 3x+1 on one side, to get rid of that silly fraction.

OpenStudy (anonymous):

3x+1=20, then?

OpenStudy (anonymous):

Great job! then what, you tell me.

OpenStudy (anonymous):

subtract one from both sides and divide by three

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes! so what is x?

OpenStudy (anonymous):

Well If I didn't have options, I would attempt to make a decimal; however my options are x=19/3, x=7, x=3/4, and x=17... so x=19/3? :)

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

Thank you oh so much :):):)

OpenStudy (anonymous):

\[\frac{ 3x }{ 4 }+\frac{ 1 }{ 4 }=5\] \[\frac{ 3x+1 }{ 4 }=5\] \[3x+1=20\] \[3x=19\] \[x=\frac{ 19 }{ }\] do you understand everything

OpenStudy (anonymous):

you know, there is also another method of solving this... it is called plugging in.

OpenStudy (anonymous):

ukk x=19/3 as my final

OpenStudy (anonymous):

Haha, thank you both, I do understand now. I feel so silly because of how simple it was :p

OpenStudy (anonymous):

you're welcome :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

and if you ever get stuck , you can always plug in each value that x might equal to get 5... until you get the right answer.

OpenStudy (anonymous):

Thank you, also very helpful :):)

OpenStudy (anonymous):

so in this case, \[\frac{ 3(\frac{ 19 }{ 3 }) }{ 4 }=5\]

OpenStudy (anonymous):

and you're welcome ☺